Two kilograms of hamburger and three kilograms of chicken cost $5.55. One kilogram of hamburger and two kilograms of chicken cost $3.37. Find the cost of one kilogram of hamburger and the cost of kilogram of chicken?
Found 2 solutions by mananth, rwm:
Answer by mananth(16949)

Hamburger------------x cost
chicken ------------ y cost
2x+3y=5.55 .............1
x+2y=3.37 .............2
Eliminate y
multiply (1)by -2
Multiply (2) by 3
-4x-6y=-11.1
3x+6y=10.11
Add the two equations
-1x=-0.99
x = 0.99
plug value of x in (1)
2x+3y=5.55
1.98 +3y=5.55
3y=5.55 -1.98
3y=3.57
y=1.19

Answer by rwm(914)

2h+3c=555
1h+2c=337
h=99
c=119